OpenStudy (anonymous):

determine k so that 4x^2-3x+k has one root which is twice the other root..

OpenStudy (anonymous):

help

OpenStudy (anonymous):

(4x - a)(x - b) 4x^2 -(4b+a)x + ab 1. 4b + a = 3 2. ab = k 3. a = 2b sub 3 into 1: 4b + 2b = 3 b = 1/2 sub that into 3: a = 1 and into 2: k = 1/2 the graph: http://bit.ly/pqxMPt

OpenStudy (anonymous):

:)